OpenStudy (dls):

Why does NH4NO3 dissolve in water spontaneoulsy even when this process is endothermic?

OpenStudy (shubhamsrg):

dG = dH - TdS

OpenStudy (dls):

:)

OpenStudy (shubhamsrg):

:)

OpenStudy (dls):

so what

OpenStudy (shubhamsrg):

dH>0, but still dG<0 since TdS must be >dH

OpenStudy (dls):

so what

OpenStudy (shubhamsrg):

so lets dance :)

OpenStudy (dls):

i wrote the wrong answer -.-

OpenStudy (shubhamsrg):

ohh, how was the exam ?

OpenStudy (dls):

hadd se zda bekar

OpenStudy (shubhamsrg):

hmm, matlab 100/100 nahi ayenge, :/

OpenStudy (dls):

I wrote being endothermic reaction isn't the only criterion for spontaneity

OpenStudy (shubhamsrg):

98 ya 99

OpenStudy (shubhamsrg):

thik to hai

OpenStudy (dls):

ice melts,diffusion,etc all are exothermic

OpenStudy (dls):

not diffusion is exo i mean i wrote only THAT mucb

OpenStudy (shubhamsrg):

mera comment niche kaise aa gay ? o.O

OpenStudy (dls):

much*

OpenStudy (shubhamsrg):

gaya*

OpenStudy (dls):

idk

OpenStudy (shubhamsrg):

well, its never too late to dance :)

OpenStudy (dls):

is it wrong?

OpenStudy (shubhamsrg):

1/2 no. to mil hi jayega

OpenStudy (dls):

wtf

OpenStudy (dls):

0 ya 1 deti hai wo

OpenStudy (shubhamsrg):

hmm. binary digits ?

OpenStudy (dls):

write chemical test to distinguish b/w but 1 yne aur but 2 yne

OpenStudy (shubhamsrg):

ammonical AgNo3

OpenStudy (dls):

:O?

OpenStudy (dls):

i wrote HBR o_O

OpenStudy (shubhamsrg):

1-yne ka test hota hai amm. AgNO3ya for amm. CuCl

OpenStudy (shubhamsrg):

fir*

OpenStudy (dls):

HBR is wrong? :P

OpenStudy (shubhamsrg):

nahi, thats the trick, HBr is never wrong :)

OpenStudy (dls):

wtf

OpenStudy (dls):

HBr se react karne pe one gives but 1 bromo butane and other 2

OpenStudy (shubhamsrg):

tum dekh ke kaise pata karoge ki 1 bromo bana hai ki 2bromo ?

OpenStudy (shubhamsrg):

amm wala test color change hota hai.

OpenStudy (shubhamsrg):

wahi sahi ans hai

OpenStudy (dls):

acha

OpenStudy (dls):

ok

OpenStudy (shubhamsrg):

and, HBr , is never wrong anyways :)

OpenStudy (dls):

why?

OpenStudy (shubhamsrg):

because YOU wrote it :)

OpenStudy (dls):

The pH of a monobasic acid HA is 4.50.Calculate concentration of everything and Kc.

OpenStudy (dls):

You did it=you are god :P

OpenStudy (dls):

spent 30 minutes and nothing

OpenStudy (dls):

[H+]=[A-]=3.33*10^-5 thats all i got

OpenStudy (shubhamsrg):

hmm..

OpenStudy (shubhamsrg):

HA - > H+ + A- 1 0 0 1-a a a a= 3,33*10^-5 1-a = ? Kc = a^2 / 1-a = a^2 approximately

OpenStudy (dls):

how can u take it `1

OpenStudy (shubhamsrg):

unit conc. the way it is done B|

OpenStudy (dls):

chal na aise he kuch bhi